OpenStudy (anonymous):

for the function f(x) = x + 5 what is the ordered pair for the point on the graph when x = 4w (4w x5) (4w 4w + 5) (x x + 5) (x 4w + 5) yahoo

OpenStudy (anonymous):

(4w ; 4w + 5) You just have to substitute on the f(x) like f(x) = x + 5 f(4w) = 4w +5 which means (4w ; 4w+5) what means when X = 4w, Y = 4w + 5

OpenStudy (anonymous):

thks bro got it